# An extension of Birkhoff’s representation theorem to locally-finite distributive lattices

### 06 Mar 2026 v1

@@ 1. Introduction
In combinatorics many interesting lattices aren’t finitary but they are locally-finite, so it is desirable to extend the representation theorem to locally-finite lattices in some way. Two obstacles immediately present themselves. One is that some locally-finite lattices have no join-irreducible elements at all, making it unclear what a representation could be constructed from. The premier example of this is $\mathbb{Z}\times\mathbb{Z}$, as shown in fig. 2.
@@
The second obstacle is that in many useful cases, the lattice is isomorphic to the lattice of some, but not all, of the order ideals of join-irreducible elements, all of which are infinite. Often these ideals consist of all the ideals with the exception of the empty ideal and the ideal of all join-irreducible elements. A typical example is shown in fig. 3.
@@
Thus our major goal is a general theory of representation of locally-finite distributive lattices with a secondary goal of understanding lattices that are isomorphic to the lattice of some, but not all, order ideals of its join-irreducible elements.
@@ 2. Representation of non-finitary lattices
Our analysis of general distributive lattices follows the discussion in [Nat2017a] and our references to it refer to chapter 8 if not otherwise specified. [^1] Prof. Nation informs me that the discussion in [Nat2017a] is derived from Bob Dilworth’s lattice theory course at Caltech in 1970–71.
This discussion is an alternative to Stone’s presentation of a representation [Stone1938a] that avoids using the machinery of topology, which makes its application to discrete lattices clearer.
@@
Let $P$ be a poset. We define that $F\subset P$ is an order filter (of $P$) iff $x\in P$ and $x\leq z$ imply $z\in F$. We define that $I\subset P$ is an order ideal (of $P$) iff $x\in P$ and $z\leq x$ imply $z\in I$.
@@
[Nat2017a]*sec. 7 We define that $f\subset\mathcal{L}$ is a lattice filter (of $\mathcal{L}$) iff (1) $f$ is non-empty, (2) $x,y\in f$ implies $x\vee y\in f$, and (3) $x\in f$ and $x\leq z$ imply $z\in f$. We define that $i\subset\mathcal{L}$ is a lattice ideal (of $\mathcal{L}$) iff (1) $i$ is non-empty, (2) $x,y\in i$ implies $x\wedge y\in i$, and (3) $x\in i$ and $z\leq x$ imply $z\in i$.
@@
[Nat2017a] We define that $f$ is a proper lattice filter (of $P$) iff $f$ is a lattice filter and $f$ is not $\mathcal{L}$. We define a proper lattice ideal dually.
@@
We define $\mathcal{F}$ to be the poset of lattice filters of $\mathcal{L}$, ordered by inverse inclusion.
For $\mathcal{L}=\mathbb{Z}\times\mathbb{Z}$, some proper filters are shown in fig. 2 and the corresponding lattice $\mathcal{F}$ is shown in fig. 5.
Figure 4. The lattice of filters $\mathcal{F}$ for $\mathcal{L}=\mathbb{Z}\times\mathbb{Z}$ with its prime elements outlined, all of which are sur-primes
@@
Given $f,g\in\mathcal{F}$, we define $f\mathop{\wedge\mkern-8.0mu\wedge}g=\{x\wedge y\,|\,x\in f\textup{ and }y\in g\}$. [^2] We pronounce $f\mathop{\wedge\mkern-8.0mu\wedge}g$ as “f union-meet g” due to lem. 9.
@@
Given $\in\mathcal{F}, fqf\mathop{\wedge\mkern-8.0mu\wedge}q$ contains $f$ and $q$.
@@
$f\mathop{\wedge\mkern-8.0mu\wedge}g$ is a lattice filter and is the meet-closure of $f\cup g$.
@@
Regarding $f\mathop{\wedge\mkern-8.0mu\wedge}g$ is a lattice filter: Given $x\in f$, $y\in g$ and $z\in\mathcal{L}$ with $x\wedge y\leq z$ then $z\vee x\in f$ and $z\vee y\in g$. By distributivity, $(z\vee x)\wedge(z\vee y)=z\vee(x\wedge y)=z$, showing $z\in f\mathop{\wedge\mkern-8.0mu\wedge}g$.
Regarding $f\mathop{\wedge\mkern-8.0mu\wedge}g$ is contained in the meet-closure of $f\cup g$: Trivially, $f\mathop{\wedge\mkern-8.0mu\wedge}g$ is contained in the meet-closure of $f\cup g$.
Regarding $f\mathop{\wedge\mkern-8.0mu\wedge}g$ contains meet-closure of $f\cup g$: Since meet is commutative, any element $z$ in the meet-closure of $f\cup g$ must be the meet of finite non-empty subsets of elements of $f$ and $g$, $z=x_{1}\wedge x_{2}\wedge x_{3}\wedge\cdots\wedge x_{n}\wedge y_{1}\wedge y_{2}\wedge y_{3}\wedge\cdots\wedge y_{m}$ with $n$ and $m$ not both 0, $x_{i}\in f$, and $y_{i}\in g$. If $n=0$, then $z\in g$, and by lem. 8, $z\in f\mathop{\wedge\mkern-8.0mu\wedge}g$. Similarly, if $m=0$, then $z\in f$, and so $z\in f\mathop{\wedge\mkern-8.0mu\wedge}g$. Otherwise, since $f$ and $g$ are meet-closed themselves, $x^{\prime}=x_{1}\wedge x_{2}\wedge x_{3}\wedge\cdots\wedge x_{n}\in f$ and $y^{\prime}=y_{1}\wedge y_{2}\wedge y_{3}\wedge\cdots\wedge y_{m}\in g$, so $z=x^{\prime}\wedge y^{\prime}\in f\mathop{\wedge\mkern-8.0mu\wedge}g$. ∎
@@
$\mathcal{F}$ is a distributive lattice.
@@
Since we know $\mathcal{F}$ is a lattice, it suffices to prove that for any $f,g,h\in\mathcal{F}$, $f\cap(g\mathop{\wedge\mkern-8.0mu\wedge}h)=(f\cap g)\mathop{\wedge\mkern-8.0mu\wedge}(f\cap h)$, which is equivalent to showing that for any $x\in\mathcal{L}$, $x\in f\cap(g\mathop{\wedge\mkern-8.0mu\wedge}h)$ iff $x\in(f\cap g)\mathop{\wedge\mkern-8.0mu\wedge}(f\cap h)$. The following are equivalent:
@@
Then choose $y^{\prime}=x\vee y$ and $z^{\prime}=x\vee z$. To derive (1) from (2): Choose $y=y^{\prime}$ and $z=z^{\prime}$, and since $f$ is closed under meets. Continuing both cases, (1) is equivalent to:
@@
[Nat2017a] A lattice filter $f$ is prime iff $\mathcal{L}\setminus f$ is a lattice ideal iff $\mathcal{L}\setminus f$ is a prime lattice ideal. And dually for lattice ideals.
Theorem 14.
[Nat2017a] For $\in\mathcal{L}, x{{x}\mathclose{\uparrow}}$ is a prime lattice filter iff $x$ is join-irreducible in $\mathcal{L}$.
Note that if $\mathcal{L}$ is not distributive, the condition in this theorem must be modified to “$x$ is join-prime”.
@@
We define $\mathcal{P}$ to be the poset of prime elements of $\mathcal{F}$, with the order (inverse inclusion) inherited from $\mathcal{F}$.
@@
For $x\in\mathcal{L}$, we define ${{x}\mathclose{\uparrow}}=\{y\in\mathcal{L}\,|\,x\leq y\}$. For $x\in\mathcal{L}$, we define ${{x}\mathclose{\downarrow}}=\{y\in\mathcal{L}\,|\,y\leq x\}$.
@@
    ${{x}\mathclose{\uparrow}}\subset p$, and
@@
    $x\in p$.
Theorem 18.
@@
In light of th. 14 we can consider $\mathcal{F}$ to be an extension of $\mathcal{L}$ which “adds the missing join-irreducible elements of $\mathcal{L}$.
Theorem 19.
@@
Define $\{h\in\mathcal{F}\,|\,h\leq f\textup{ and }h\not\leq{{x}\mathclose{\uparrow}}\}. HH$ is non-empty because $f\in H$. $H$ is a poset ordered by inverse inclusion, inherited from $\mathcal{F}$. Using the Hausdorff maximal principle, select $P$ which is a maximal totally ordered subset of $H$. Since $H$ is non-empty, $P$ is non-empty. Define $p=\bigcup P$. Because $P$ is a non-empty nested set of lattice filters, $\bigcup P$ is a lattice filter also, so $p\in\mathcal{F}$.
Because $x$ is in no member of $P$, $x\not\in\bigcup P=p$ and $p\not\leq{{x}\mathclose{\uparrow}}$.
Since $P$ contains at least one element $q$, $q\leq f$, $f\subset q$, $f\subset\bigcup P=p$, so $p\leq f$.
@@
An element of $\mathcal{P}$ (which must be prime in $\mathcal{F}$) but is not ${x}\mathclose{\uparrow}$} for any $x\in\mathcal{L}$ (which would have to be join-irreducible) we define to be a sur-prime or sur-irreducible of $\mathcal{F}$ (or of $\mathcal{P}$ or $\mathcal{L}$, by abuse of language).
For $\mathcal{L}=\mathbb{Z}\times\mathbb{Z}$, fig. 5 shows that all primes are sur-primes, and not ${x}\mathclose{\uparrow}$} for $x\in\mathcal{L}$. Conversely, added elements need not be sur-primes: $\mathbb{Z}\times\mathbb{Z}$ is in $\mathcal{F}$, not a ${x}\mathclose{\uparrow}$}, but also is not prime.
Theorem 22.
$f\in\mathcal{F}$ is join-irreducible in $\mathcal{F}$ iff $f$ is a prime lattice filter.
@@
Regarding $\Rightarrow$: Assume $f$ is join-reducible. Since $f$ is join-reducible, there exist filters $g$ and $h$ for which $g,h\neq f$ and $f=g\vee h=g\cap h$. Choose $x\in g\setminus f$ and $y\in h\setminus f$; thus $x,y\not\in f$. Since $g$ and $h$ are filters, $x\vee y\in g,h$, and so $x\vee y\in f$. Thus $f$ is not prime.
Regarding $\Leftarrow$: Assuming $f$ is not prime, there exists $x$ and $y$ with $x,y\not\in f$ but $x\vee y\in f$. Define $g={{x}\mathclose{\uparrow}}\mathop{\wedge\mkern-8.0mu\wedge}f$ and $h={{y}\mathclose{\uparrow}}\mathop{\wedge\mkern-8.0mu\wedge}h$, both of which are $\neq f$. We will show that $f=g\vee h=g\cap f$. Clearly $f\subset g\cap h$. Conversely, suppose $z\in g\cap h$. Then there exists $u,u^{\prime},v,v^{\prime}\in f$ for which $z=u\wedge u^{\prime}=v\wedge v^{\prime}$, $u\geq x$, $v\geq y$, and $u^{\prime},v^{\prime}\in f$. Then $z=(u\wedge u^{\prime})\vee(v\wedge v^{\prime})=(u\vee v)\wedge(u\vee v^{\prime})\wedge(u^{\prime}\vee v)\wedge(u^{\prime}\vee v^{\prime})$. But all four terms of that meet are $\in f$, so $z\in f$, showing $g\cap h\subset f$. ∎
@@
We define $\mathcal{I}$ to be the lattice of order ideals of $\mathcal{P}$ under inclusion, which has $\cap$ as meet and $\cup$ as join.
@@
Application to finitary lattices

### 15 Mar 2026 v2

@@ 1. Introduction
In combinatorics many interesting lattices aren’t finitary but they are locally-finite, so it is desirable to extend the representation theorem to locally-finite lattices in some way. Two obstacles immediately present themselves. One is that some locally-finite lattices have no join-irreducible elements at all, making it unclear what a representation could be constructed from. A premier example of this is $\mathbb{Z}\times\mathbb{Z}$, as shown in fig. 2.
@@
The second obstacle is that in many useful cases, the lattice is isomorphic to the lattice of some, but not all, of the order ideals of join-irreducible elements. Often these ideals consist of all the ideals with the exception of the empty ideal and the ideal of all join-irreducible elements. A typical example is shown in fig. 3.
@@
Thus our major goal is a general theory of representation of locally-finite distributive lattices. A secondary goal is understanding lattices that are isomorphic to the lattice of some, but not all, order ideals of their join-irreducible elements.
@@ 2. Representation of non-finitary lattices
First we will lay out the analysis of general distributive lattices. Our analysis follows the discussion in [Nat2017a] and our references to it refer to chapter 8 if not otherwise specified. [^1] Prof. Nation informs me that the discussion in [Nat2017a] is derived from Bob Dilworth’s lattice theory course at Caltech in 1970–71. This discussion is an alternative to Stone’s presentation of a representation [Stone1938a] that avoids using the machinery of topology, which makes its application to discrete lattices clearer. Following the general analysis, we will apply it to locally-finite lattices.
@@
Let $P$ be a poset. We define that $F\subset P$ is an order filter (of $P$) iff $x\in F$ and $x\leq z$ imply $z\in F$. We define that $I\subset P$ is an order ideal (of $P$) iff $x\in F$ and $z\leq x$ imply $z\in I$.
@@
[Nat2017a]*sec. 7 Let $\mathcal{L}$ be a lattice. We define that $f\subset\mathcal{L}$ is a lattice filter (of $\mathcal{L}$) iff (1) $f$ is non-empty, (2) $x,y\in f$ implies $x\wedge y\in f$, and (3) $x\in f$ and $x\leq z$ imply $z\in f$. We define that $i\subset\mathcal{L}$ is a lattice ideal (of $\mathcal{L}$) iff (1) $i$ is non-empty, (2) $x,y\in i$ implies $x\vee y\in i$, and (3) $x\in i$ and $z\leq x$ imply $z\in i$.
@@
[Nat2017a] We define that $f$ is a proper lattice filter (of a lattice $\mathcal{L}$) iff $f$ is a lattice filter and $f$ is $\neq\mathcal{L}$. We define a proper lattice ideal dually.
There is of course a natural duality on lattices, so for many facts we prove, we also state the dual fact without giving the dual proof.
@@
We define $\mathcal{F}$ to be the poset of lattice filters of $\mathcal{L}$, ordered by inverse inclusion. Dually, we define $\mathcal{F}^{\prime}$ to be the poset of lattice ideals of $\mathcal{L}$, ordered by inclusion.
For $\mathcal{L}=\mathbb{Z}\times\mathbb{Z}$, some filters (all proper) are shown in fig. 2 and the corresponding lattice $\mathcal{F}$ is shown in fig. 5.
Figure 4. The lattice of filters $\mathcal{F}$ for $\mathcal{L}=\mathbb{Z}\times\mathbb{Z}$ with its prime elements outlined, all of which are secondary primes
@@
Given $f,g\in\mathcal{F}$, we define $f\mathop{\wedge\mkern-8.0mu\wedge}g=\{x\wedge y\,|\,x\in f\textup{ and }y\in g\}$. Dually, given $f^{\prime},g^{\prime}\in\mathcal{F}^{\prime}$, we define $f^{\prime}\mathop{\vee\mkern-8.0mu\vee}g^{\prime}=\{x\vee y\,|\,x\in f^{\prime}\textup{ and }y\in g^{\prime}\}$. [^2] We pronounce $f\mathop{\wedge\mkern-8.0mu\wedge}g$ as “$f$ union-meet gg” due to lem. 9. Dually, we pronounce $f^{\prime}\mathop{\vee\mkern-8.0mu\vee}g^{\prime}$ as $f^{\prime}$ union-join $g^{\prime}$”.
@@
Given $\in\mathcal{F}, fgf\mathop{\wedge\mkern-8.0mu\wedge}g$ contains $f$ and $g$. Dually, given $f^{\prime},g^{\prime}\in\mathcal{F}^{\prime}$, $f^{\prime}\mathop{\vee\mkern-8.0mu\vee}g^{\prime}$ contains $f^{\prime}$ and $g^{\prime}$.
@@
Given $\in\mathcal{F}, fgf\mathop{\wedge\mkern-8.0mu\wedge}g$ is a lattice filter and is the meet-closure of $f\cup g$. Dually, given $f^{\prime},g^{\prime}\in\mathcal{F}^{\prime}$, $f^{\prime}\mathop{\wedge\mkern-8.0mu\wedge}g^{\prime}$ is a lattice ideal and is the join-closure of $f^{\prime}\cup g^{\prime}$.
@@
Regarding $f\mathop{\wedge\mkern-8.0mu\wedge}g$ is closed upward: Given $x\in f$, $y\in g$ and $z\in\mathcal{L}$ with $x\wedge y\leq z$ then $z\vee x\in f$ and $z\vee y\in g$. By distributivity, $(z\vee x)\wedge(z\vee y)=z\vee(x\wedge y)=z$, showing $z\in f\mathop{\wedge\mkern-8.0mu\wedge}g$.
Regarding $f\mathop{\wedge\mkern-8.0mu\wedge}g$ is contained in the meet-closure of $f\cup g$: Trivial by def. 7.
Regarding $f\mathop{\wedge\mkern-8.0mu\wedge}g$ contains the meet-closure of $f\cup g$: This follows from lem. 8 and the fact that $f\mathop{\wedge\mkern-8.0mu\wedge}g$ is meet-closed. ∎
@@
$\mathcal{F}$ is a distributive lattice. Dually, $\mathcal{F}^{\prime}$ is a distributive lattice.
@@
Since we know $\mathcal{F}$ is a lattice, it suffices to prove that for any $f,g,h\in\mathcal{F}$, $f\cap(g\mathop{\wedge\mkern-8.0mu\wedge}h)=(f\cap g)\mathop{\wedge\mkern-8.0mu\wedge}(f\cap h)$, which is equivalent to showing that for any $x\in\mathcal{L}$, $x\in f\cap(g\mathop{\wedge\mkern-8.0mu\wedge}h)$ iff $x\in(f\cap g)\mathop{\wedge\mkern-8.0mu\wedge}(f\cap h)$. To show that, the following are equivalent:
@@
Then choose $y^{\prime}=x\vee y$ and $z^{\prime}=x\vee z$. To derive (1) from (2): Choose $y=y^{\prime}$ and $z=z^{\prime}$, and note that $f$ is closed under meets. Thus, (1) is equivalent to:
@@
[Nat2017a] A lattice filter $f$ is prime iff $\mathcal{L}\setminus f$ is a lattice ideal iff $\mathcal{L}\setminus f$ is a prime lattice ideal. Dually, a lattice ideal $f$ is prime iff $\mathcal{L}\setminus f$ is a lattice filter iff $\mathcal{L}\setminus f$ is a prime lattice filter.
@@
For $x\in\mathcal{L}$, we define ${{x}\mathclose{\uparrow}}=\{y\in\mathcal{L}\,|\,x\leq y\}$, the principal filter generated by $x$. For $x\in\mathcal{L}$, we define ${{x}\mathclose{\downarrow}}=\{y\in\mathcal{L}\,|\,y\leq x\}$, the principal ideal generated by $x$.
Lemma 15.
[Nat2017a] For $\in\mathcal{L}, x{{x}\mathclose{\uparrow}}$ is a prime lattice filter iff $x$ is join-irreducible in $\mathcal{L}$. Dually, for $\in\mathcal{L}, x{{x}\mathclose{\downarrow}}$ is a prime lattice ideal iff $x$ is meet-irreducible in $\mathcal{L}$.
Note that if $\mathcal{L}$ were not distributive, the conditions in this theorem would have to be modified to “$x$ is join-prime” and “$x$ is meet-prime”.
@@
We define $\mathcal{P}$ to be the poset of prime elements of $\mathcal{F}$, with the order (inverse inclusion) inherited from $\mathcal{F}$. Dually, we define $\mathcal{P}^{\prime}$ to be the poset of prime elements of $\mathcal{F}^{\prime}$, with the order (inclusion) inherited from $\mathcal{F}^{\prime}$.
@@
Lem. 13 shows that the $\omega:\mathcal{P}\rightarrow\mathcal{P}^{\prime}:p\mapsto\mathcal{L}\setminus p$ is a poset isomorphism between $\mathcal{P}$ and $\mathcal{P}^{\prime}$.
Lemma 18.
Given $x\in\mathcal{L}$ and $f\in\mathcal{F}$, the following are equivalent:
@@
    ${{x}\mathclose{\uparrow}}\subset f$, and
@@
    $x\in f$.
Dually, given $x\in\mathcal{L}$ and $f^{\prime}\in\mathcal{F}^{\prime}$, the following are equivalent:
  1. (1)
    $f^{\prime}\leq{{x}\mathclose{\downarrow}}$,
  2. (2)
    $f^{\prime}\subset{{x}\mathclose{\downarrow}}$, and
  3. (3)
    $x\in f^{\prime}$.
Lemma 19.
@@
In light of lem. 15 we can consider $\mathcal{F}$ to be an extension of $\mathcal{L}$ which “adds the missing join-irreducible elements of $\mathcal{L}$.”
Lemma 20.
@@
Define $\{h\in\mathcal{F}\,|\,h\leq f\textup{ and }h\not\leq{{x}\mathclose{\uparrow}}\}. HH$ is non-empty because $f\in H$. $H$ is a poset ordered by inverse inclusion, inherited from $\mathcal{F}$. Using the Hausdorff maximal principle, select $P$ which is a maximal totally ordered subset of $H$. Since $H$ is non-empty, $P$ is non-empty. Define $p=\bigcup P$. Because $P$ is a non-empty nested set of lattice filters, $\bigcup P$ is also a lattice filter, so $p\in\mathcal{F}$.
Because $x$ is in no member of $P$, $x\not\in\bigcup P=p$ and so $p\not\leq{{x}\mathclose{\uparrow}}$.
Since $P$ contains at least one element $q$, $q\leq f$, $f\subset q$, and $f\subset\bigcup P=p$, so $p\leq f$.
@@
Let $p$ be an element of $\mathcal{P}$ (which thus is prime in $\mathcal{F}$). If $p={{x}\mathclose{\uparrow}}$ for some $x\in\mathcal{L}$ (which would have to be join-irreducible) we define $p$ to be a principal prime. If there is no such $x$, then we define $p$ to be a secondary prime. Both terms are relative to $\mathcal{L}$, to $\mathcal{F}$, or to $\mathcal{P}$ (by abuse of language).
For $\mathcal{L}=\mathbb{Z}\times\mathbb{Z}$, fig. 5 shows that all primes are secondary primes. Conversely, there can be elements of $\mathcal{F}$ that are not of the form ${x}\mathclose{\uparrow}$} and are not secondary primes: $\mathbb{Z}\times\mathbb{Z}$ is in $\mathcal{F}$, is not an ${x}\mathclose{\uparrow}$}, but also is not prime.
Lemma 23.
$f\in\mathcal{F}$ is join-irreducible in $\mathcal{F}$ iff $f$ is a prime lattice filter. Dually, $f^{\prime}\in\mathcal{F}^{\prime}$ is meet-irreducible in $\mathcal{F}^{\prime}$ iff $f$ is a prime lattice ideal.
@@
Regarding $\Rightarrow$: Assume $f$ is join-reducible. Then there exist filters $g$ and $h$ for which $g,h\neq f$ and $f=g\vee h=g\cap h$. Choose $x\in g\setminus f$ and $y\in h\setminus f$; thus $x,y\not\in f$. Since $g$ and $h$ are filters, $x\vee y\in g,h$, and so $x\vee y\in f$. Thus $f$ is not prime.
Regarding $\Leftarrow$: Assume $f$ is not prime. Then there exists $x$ and $y$ with $x,y\not\in f$ but $x\vee y\in f$. Define $g={{x}\mathclose{\uparrow}}\mathop{\wedge\mkern-8.0mu\wedge}f$ and $h={{y}\mathclose{\uparrow}}\mathop{\wedge\mkern-8.0mu\wedge}f$, both of which are $\neq f$. We will show that $f=g\vee h=g\cap f$. By lem. 8, $f\subset g\cap h$. Conversely, suppose $z\in g\cap h$. Then there exists $u,u^{\prime},v,v^{\prime}\in f$ for which $z=u\wedge u^{\prime}=v\wedge v^{\prime}$, $u\geq x$, $v\geq y$, and $u^{\prime},v^{\prime}\in f$. Then $z=(u\wedge u^{\prime})\vee(v\wedge v^{\prime})=(u\vee v)\wedge(u\vee v^{\prime})\wedge(u^{\prime}\vee v)\wedge(u^{\prime}\vee v^{\prime})$. But all four terms of that meet are $\in f$, so $z\in f$, showing $g\cap h\subset f$. Thus, $f=g\cap h=g\vee h$, showing $f$ is join-reducible. ∎
@@
We define $\mathcal{I}$ to be the lattice of order ideals of $\mathcal{P}$ under inclusion, which has $\cap$ as meet and $\cup$ as join. Dually, $\mathcal{I}$ is lattice-isomorphic to the lattice of order ideals of $\mathcal{P}^{\prime}$.
@@
3. Application to finitary lattices
Applying the general theory to finitary lattices is straightforward.